• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name aniol

    Log entry time 20:24:46 on June14,2004

    Entry number 125759

    keyword=septa hung up in r2131

    Brian noticed that the spectrometer detectors were not showing electrons, i.e., only pedestals. We found that the pan detector guis agreed with this, i.e., there were no electrons in the spectrometers.
    Looking at the septa control screens we noticed that all the readbacks were constant in both septa. We called Heidi who walked us through a reboot of the septa PLC control and ioc controls over the telephone. Unfortunately, this did not give us control over the septa again. She is on her way to the lab.